# OccScene: Semantic Occupancy-based Cross-task Mutual Learning for 3D Scene Generation

### v1

@@ IV-D Ablation Study
Joint Diffusion Scheme (JDS). As shown in Table VI, to conduct the setting of ‘w/o JDS’ for effect evaluation, we detach the perception model and predict the semantic occupancy in an offline manner. Specifically, the generative framework is conditioned on the pre-produced occupancy to generate corresponding images, and the generated images of the last inference step are leveraged to train the perception model. As illustrated in the table, the joint diffusion scheme benefits the fidelity of image and video generation significantly. Moreover, the joint scheme enhances 2.76 mIoU for the perception performance compared to the offline strategy, which stems from the utilization of different information capacities in the generation process.
@@
TABLE VII: Comparison of generation quality with different architecture designs of the Mamba-based Dual Alignment (MDA) module on the SemanticKITTI validation set.
Mamba-based Dual Alignment. As shown in Figure 10, to evaluate the effect of the MDA module, we design and compare different encoding architectures including Attention-based encoding, GRU-based encoding and Mamba-based encoding. Note that the setting of ‘w/o MDA’ in Table VI is conducted with cross-attention without Cross-view Camera Encoding. Please refer to the supplementary material for architectural details of different designs.
@@
As shown in Table VII, the mamba-based demonstrates superior running speed and generation quality, which we attribute to the linear-complexity operator and efficient long-term modeling to process input high-dimensional data in a single pass. Specifically, the mamba-based architecture reduces 32.52% running time compared to the attention-based design and 19.07% FID compared to the GRU-based design, respectively.
@@
TABLE VIII: FID scores of the proposed method with different sampling steps. The evaluations are conducted on the NYUv2 test set, SemanticKITTI validation set and Nuscene-Occupancy validation set, respectively.
Efficiency Analyse. We report the running time and generation quality of several schemes across different datasets with our proposed OccScene on the NVIDIA A100 GPU, which are detailed in Table VIII. It’s worth noting that our method could effectively achieve compelling performance gains with acceptable time consumption. The results reveal that beyond 50 sampling steps, the marginal gains in effectiveness are minimal relative to the increased computational time. Consequently, we have selected 50 sampling steps as the default configuration, which provides an optimal balance between efficiency and effectiveness.

### v2

@@ IV-D Ablation Study
Joint Diffusion Scheme (JDS). As shown in Table VII, to conduct the setting of ‘w/o JDS’ for effect evaluation, we detach the perception model and predict the semantic occupancy in an offline manner. Specifically, the generative framework is conditioned on the pre-produced occupancy to generate corresponding images, and the generated images of the last inference step are leveraged to train the perception model. As illustrated in the table, the joint diffusion scheme benefits the fidelity of image and video generation significantly. Moreover, the joint scheme enhances 2.04 mIoU for the perception performance compared to the offline strategy, which stems from the utilization of different information capacities in the generation process.
[…]
MonoScene [23] 42.51 26.94
MonoScene [23] + OccScene (detached gradients) 43.05 27.47
MonoScene [23] + OccScene (attached gradients) 44.34 29.78
NDC-Scene [83] 44.17 29.03
NDC-Scene [83] + OccScene (detached gradients) 45.62 30.20
NDC-Scene [83] + OccScene (attached gradients) 47.89 32.12
ISO [84] 47.11 31.25
ISO [84] + OccScene (detached gradients) 48.60 32.09
ISO [84] + OccScene (attached gradients) 50.07 33.92
TABLE VIII: Effect of the joint training methodology on the NYUv2 test set, which effectively improves extensive perception models.
We evaluate the effect of the joint training methodology with different perception models on the NYUv2 test set, as shown in Table VIII. Specifically, the setting of ‘OccScene (detached gradients)’ represents training baseline models (e.g., MonoScene [23], NDC-Scene [83], ISO [84]) with gradients detached from the generative model but with identical data augmentation (including augmented noisy data). The setting of ‘OccScene (attached gradients)’ represents training baseline models with gradients attached to the generative model. The ‘attached gradients’ setting consistently outperforms the ‘detached gradients’ setting across all baseline models, improving 3.72 IoU and 3.09 mIoU for NDC-Scene [83]. This performance enhancement underscores the effectiveness of our joint training methodology in improving extensive perception models.
@@
TABLE IX: Comparison of generation quality with different architecture designs of the Mamba-based Dual Alignment (MDA) module on the SemanticKITTI validation set.
Mamba-based Dual Alignment. As shown in Figure 11, to evaluate the effect of the MDA module, we design and compare different encoding architectures including Attention-based encoding, GRU-based encoding and Mamba-based encoding. Note that the setting of ‘w/o MDA’ in Table VII is conducted with cross-attention without Cross-view Camera Encoding. Please refer to the supplementary material for architectural details of different designs.
@@
As shown in Table IX, the mamba-based demonstrates superior running speed and generation quality, which we attribute to the linear-complexity operator and efficient long-term modeling to process input high-dimensional data in a single pass. Specifically, the mamba-based architecture reduces 32.52% running time compared to the attention-based design and 19.07% FID compared to the GRU-based design, respectively.
[…]
w/o MDA-D 22.12 121.84 14.36
w/o MDA-T 23.02 128.17 14.04
w/ MDA 19.86 113.28 14.98
TABLE X: Ablation study of the Mamba-based Dual Alignment (MDA) module applied along different dimensions on the SemanticKITTI validation set.
Table X illustrates the effect of applying the MDA module along different dimensions. For the setting of ‘ w/o MDA-D’, the bidirectional Mamba block is applied exclusively to the video diffusion latent $\bm{L}\in C_{L}\times N\times H_{L}\times W_{L}$, while 3D convolution layers are used for the semantic occupancy $\bm{{X}}_{occ}\in 1\times D\times H_{occ}\times W_{occ}$. Conversely, the setting of ‘ w/o MDA-D’ employs the bidirectional Mamba block only for the semantic occupancy $\bm{{X}}_{occ}$, with 3D convolutions handling the video latent feature $\bm{L}$. As shown in the table, applying the MDA module along both the depth and temporal dimensions yields substantial performance enhancements, improving FID by 2.26 and 3.16, respectively.
@@
TABLE XI: FID scores of the proposed method with different sampling steps. The evaluations are conducted on the NYUv2 test set, SemanticKITTI validation set and Nuscene-Occupancy validation set, respectively.
Efficiency Analyse. We report the running time and generation quality of several schemes across different datasets with our proposed OccScene on the NVIDIA A100 GPU, which are detailed in Table XI. It’s worth noting that our method could effectively achieve compelling performance gains with acceptable time consumption. The results reveal that beyond 50 sampling steps, the marginal gains in effectiveness are minimal relative to the increased computational time. Consequently, we have selected 50 sampling steps as the default configuration, which provides an optimal balance between efficiency and effectiveness.